OpenStudy (heretohelp):

1) Jack spends 1 1/2 as long on his homework as Jill. Lask week, Jill spent 9 1/2 hours doing homework. How long did Jack spend doing homework? 2) The week before Jack spent 20 3/4 hours doing homework. How long did Jill spend doing homework? Please Help x

OpenStudy (anonymous):

Jack spends 2/3s the time Jill did Multiply 2/3 by 5 1/2 hrs 5 1/2 hrs = 5/2 So (2/3)(5/2)= 5/3 So Jack spends 5/3s hours doing his HW OR 1 and 2/3s hours if you want this in minutes multiply 5/3 by 60 and you get 100 minutes which is 1 hour and 40 minutes

OpenStudy (anonymous):

Set up an algebra expression.. (Jill)1.5=(Jack) substitute in Jill's time last week

OpenStudy (anonymous):

reversely substiute in Jack's time for 2 and solve for jill

OpenStudy (heretohelp):

Thank you both of you, you guys were a big help!